### v1

# Argumentative Debates for Transparent Bias Detection [Technical Report]

Abstract
As the use of AI in society grows, addressing emerging biases is essential to prevent systematic disadvantages against specific groups. Several fairness notions/algorithmic methods have been proposed, but, with very few exceptions, these tend to ignore transparency. Instead, interpretability and explainability are core requirements for algorithmic fairness, even more so than for other algorithmic solutions, given the human-oriented nature of fairness. In this paper, we contribute a novel transparent method for bias detection to empower debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods. Our method builds upon techniques from (formal and computational) argumentation, whereby debates result from arguing about biases within and across neighbourhoods. We evaluate our approach experimentally and demonstrate its strengths in performance against an argumentative baseline.
1 Introduction
Figure 1: A QBAF generated by our approach for the COMPAS dataset (with protected feature $X_{p}\!=\!race$ and protected group g=g\!= African-American) for two neighbourhoods of size $K=10,100$ (critical question arguments with zero strength omitted). Dashed/solid edges are supports/atttacks. Green/red edges are supports/attacks (from arguments) with nonzero strength, and black edges are (from arguments) with zero strength. Edge width reflects the strength of originating arguments. (Strengths below/above nodes. Details in Section 6.)
As the use of AI in society grows, addressing its potential unfairness is becoming increasingly crucial. For example, an unfair AI model used to support decision-making in healthcare, discriminating against a certain sub-population, would be extremely harmful. In particular, it is essential to address any potential biases against specific groups/individuals in data and AI models trained thereupon (Mehrabi et al. 2022; Caton and Haas 2024). Various notions of fairness have been proposed in the literature (Mehrabi et al. 2022; Caton and Haas 2024; Waller et al. 2024). Amongst these notions, statistical (or demographic) parity (Corbett-Davies et al. 2017) defines fairness as an equal probability of being classified with the desirable (positive) label in different groups. However, existing solutions to identify and rectify unfairness tend to focus on optimising fairness and ignore the need for transparency for the roots of (un)fairness (Caton and Haas 2024). Instead, transparency, afforded by interpretability and explainability, is crucial for the detection and mitigation of bias, as is the case for any algorithmic solution, but even more so when bias is concerned, given that fairness is inherently human-oriented (Waller et al. 2024). While considerable efforts have been made towards explainability of machine learning models (Dejl et al. 2025; Ayoobi et al. 2019, 2021, 2022, 2023a, 2023b, 2025; Leofante et al. 2024), algorithmic fairness has received limited attention (exceptions include (Grabowicz, Perello, and Mishra 2022; Waller, Rodrigues, and Cocarascu 2024)).
We propose a novel transparent method to detect bias through argumentative debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods. Debates recently emerged as a powerful mechanism to address various challenges in contemporary AI, such as AI safety (Brown-Cohen, Irving, and Piliouras 2024), and to improve LLM performance (Khan et al. 2024). In much of this literature, debates are unstructured and inform other entities that decide the debates’ outcome. Thus, while these debates provide a justification for the outcomes, these outcomes are not faithfully explained by the debates.
Our novel method for bias detection is designed to guide debates about the presence of bias against individuals. To structure debates, we rely upon argument schemes (with critical questions) (Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017), from formal argumentation. To decide the debates’ outcomes, we compute arguments’ strength with gradual semantics (Baroni, Rago, and Toni 2018) from computational argumentation. In summary:
  1. 1.
    We propose a neighbourhood-based notion of fairness for individual bias detection, adapting statistical parity.
  2. 2.
    We define novel argument schemes about the presence of bias based on (properties of) neighbourhoods.
  3. 3.
    We develop a mapping of these schemes into Quantitative Bipolar Argumentation Frameworks (QBAFs) (Baroni, Rago, and Toni 2018), enabling structured and modular reasoning for transparent bias detection (see Figure 1).
  4. 4.
    We formally connect properties of gradual semantics for QBAFs to desirable properties of bias detection.
  5. 5.
    We conduct empirical evaluations with synthetically biased models, models trained on real-world datasets, and ChatGPT-4o, consistently outperforming an argumentative baseline (Waller, Rodrigues, and Cocarascu 2024).
  6. 6.
    We show how our approach can empower debate-based bias detection in human-agent and multi-agent scenarios.
The proofs of all results in the paper can be found in Appendix B. We will release the code on acceptance.
2 Related Work
Fairness in Machine Learning.
Various methods for identifying, measuring and mitigating against bias have been proposed (Mehrabi et al. 2022; Caton and Haas 2024), but they often neglect transparency (Waller et al. 2024). One exception is (Grabowicz, Perello, and Mishra 2022), which proposed explanations to unearth and rectify bias, but relies upon feature importance, rather than debates, for transparency.
Bias can be defined in different ways, e.g. by comparing the success probability of the protected group to the one of unprotected individuals or the general population (statistical parity), or by comparing the risk of misclassification of the protected to the one of the non-protected group (predictive equality) (Corbett-Davies et al. 2017). Our approach is based on a local notion of statistical parity. That is, we identify neighbourhoods where groups are locally advantaged/disadvantaged with respect to success probability. The decision whether a bias exists is then based on the evidence provided by these neighbourhoods in a transparent way.
Debate and AI.
Some approaches rely upon debate protocols between (two) AI models trained as players in zero-sum games, e.g. towards AI safety (Irving, Christiano, and Amodei 2018; Brown-Cohen, Irving, and Piliouras 2024) or for post-hoc explaining image classifiers (Kori, Glocker, and Toni 2024). (Thauvin et al. 2024) propose a transparent model based on debate protocols between agents drawing arguments in the form of visual features. (Khan et al. 2024) use debate to resolve disagreement between (two) strong LLMs to inform and improve decisions by weaker models (LLMs or humans). In all these settings, debates are between two pre-determined entities (e.g. models or agents) and they inform decisions made by other entities (e.g. a verifier or a weaker model). Instead, in our approach, debates may be in the “mind” of a single entity or across two or more entities, as our approach focuses on providing guidance for structuring and formally evaluating (with gradual semantics) debates.
Argumentation and Bias.
(Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017) propose an argument scheme to capture arguments from bias111 //www.rationaleonline.com/map/9gy9gd/argument-from-bias, allowing to draw the conclusion that an arguer is unlikely to have taken both sides of an issue into account if the arguer is biased, subject to addressing critical questions regarding (i) evidence for the arguer’s bias and (ii) the need to take multiple sides for that issue. Our new argument scheme below is orthogonal to this and can be used in combination with it, by addressing the first critical question using properties of neighbourhoods to collect evidence for the bias of the arguer (a classifier).
(Waller, Rodrigues, and Cocarascu 2024) were the first to use gradual semantics for QBAFs to characterise bias. While we were inspired by them, our approach differs in a simpler QBAF structure (with fewer nodes and relations), the use of neighbourhoods and their properties (for more robust reasoning on bias), and the empowering of formal links between properties of semantics and of bias detection.
Argumentation and Debate.
Several have pointed to the connection between argumentation and debate, e.g. (Panisson, McBurney, and Bordini 2021; de Tarlé, Bonzon, and Maudet 2022) in general, and (Rago et al. 2025) for product recommendation. While we can leverage on these works towards debate generation, our approach is orthogonal.
3 Preliminaries
Classification Problems.
We consider variables (features) $X_{1},\dots,X_{k}$ with associated domains $D_{1},\dots,D_{k}$ and a class variable $Y$ with associated domain $\mathcal{C}$ of class labels. We use $X_{p}$ to denote a protected feature of interest and $g$ to denote its value for the protected group. Thus, $X_{p}\!=\!g$ identifies the protected group and $X_{p}\!\neq\!g$ the remaining individuals.
For $\mathcal{D}\!=\!\bigtimes_{i=1}^{k}D_{i}$, vectors $\mathbf{x}\in\mathcal{D}$ are the inputs of the classification problem. We consider (binary) classifiers $\operatorname{c}:\mathcal{D}\rightarrow\mathcal{C}\!=\!\{0,1\}$. We regard $1$ as a desirable outcome (e.g., acceptance of a loan application) and 0 as a negative outcome.
Notation 1.
Capital letters refer to variables, lowercase letters to values of variables. Bold capital letters refer to sequences of variables, bold lowercase letters to variable assignments. For example, if $\mathbf{V}\!=\!(X_{1},X_{2},X_{3})$, then $\mathbf{v}$ refers to an assignment $(x_{1},x_{2},x_{3})$ with $x_{i}\!\in\!D_{i}$ for $i\!\in\!\{1,2,3\}$. For each $\mathbf{x}\in\mathcal{D}$ and sequence of variables $\mathbf{V}=(X_{i_{1}},\dots,X_{i_{m}})$, we let $\mathbf{x}|_{\mathbf{V}}$ denote the projection of $\mathbf{x}$ onto $\mathbf{V}$. For example, if $\mathbf{x}\!=\!(a,b,c,d)$ and $\mathbf{V}\!=\!(X_{2},X_{4})$, then $\mathbf{x}|_{\mathbf{V}}\!=\!(b,d)$. Finally, we let $\mathbf{X}$ denote the sequence $(X_{1},\ldots,X_{k})$ of all variables.
QBAFs.
[…]
To assess the acceptability of arguments in a QBAF $\mathcal{Q}$, (gradual) semantics can be given by means of a strength function $\sigma_{\mathcal{Q}}:\!\mathcal{A}\!\rightarrow\![0,1]$, often defined by an iterative procedure that initializes strength values with the base scores and then repeatedly updates the strength of arguments based on the strength of their attackers and supporters. For acyclic QBAFs, this procedure always converges and is equivalent to a linear-time algorithm that first computes a topological ordering of the arguments and then updates each argument only once following the order (Potyka 2019, Proposition 3.1).
A QBAF semantics is called modular if the update function can be decomposed into an aggregation function $\operatorname{agg}(A,S)$ that aggregates the strength values $A$ of attackers and $S$ of supporters, and an influence function $\operatorname{infl}(b,a)$ that adapts the base score $b$ based on the aggregate $a$ (Mossakowski and Neuhaus 2018). Most QBAF semantics are modular. Examples include the DF-QuAD (Rago et al. 2016), Euler-based (Amgoud and Ben-Naim 2017) and quadratic energy (Potyka 2018) semantics. Aggregation functions include product $\operatorname{agg}(A,S)\!=\!\prod_{a\in A}(1-a)-\prod_{s\in S}(1-s)$ used in DF-QuAD and sum $\operatorname{agg}(A,S)\!=\!\sum_{s\in S}s-\sum_{a\in A}a$ used in Euler-based and quadratic energy semantics. The influence functions of DF-QuAD and quadratic energy both take the form $\operatorname{infl}(b,a)\!=\!b\!-\!b\cdot\!f(-s)\!+\!(1\!-b)\!\cdot f(s)$, where $f(x)\!=\!\max\{0,x\}$ for DF-QuAD and $f(x)\!=\!\frac{\max\{0,x\}}{1+\max\{0,x\}}$ for quadratic energy.
QBAF Properties.
QBAFs are often compared based on properties (Leite and Martins 2011; Amgoud and Ben-Naim 2016, 2017; Baroni, Rago, and Toni 2018), many of which are tied to properties of aggregation and influence functions (Mossakowski and Neuhaus 2018). We will use properties of (strict) monotonicity and balance of these functions from (Potyka and Booth 2024b) and recap them in Appendix A. As shown in (Potyka and Booth 2024a), product and sum aggregation satisfy balance and monotonicity, sum satisfies strict monotonicity; the influence functions of DF-QuAD and quadratic energy satisfy all properties.
4 Neighbourhoods and their Properties
The goal of this section is to make the idea of local bias precise. Intuitively, a local bias against a reference individual exists if other similar individuals are advantaged by the classifier. Formally, these similar individuals can be given by a neighbourhood of individuals surrounding the reference one.
Definition 1.
A neighbourhood is a set $\mathcal{N}\subseteq\mathcal{D}$. A neighbourhood of a point $\mathbf{x}\in\mathcal{D}$ is a set $\mathcal{N}$ such that $\mathbf{x}\in\mathcal{N}$.
If we were allowed to choose neighbourhoods arbitrarily, we may always find one that indicates presence and one that indicates absence of bias. We thus present some desirable properties of neighbourhoods. To begin with, a neighbourhood that contains only few individuals may not be reliable as they may be exceptions. The larger the neighbourhood, the larger our confidence that it is representative of the domain.
Definition 2.
Let $N\in\mathbb{N}$. A neighbourhood $\mathcal{N}\subseteq\mathcal{D}$ is called $N$-significant if $|\mathcal{N}|\geq N$.
Without further restrictions, an adversarial agent could define a neighbourhood by picking individuals systematically to demonstrate the (non-)existence of a bias. To prevent this, one reasonable assumption is that when we pick two individuals from a sample, we cannot leave out individuals between them. Mathematically, betweenness can be described by convexity:
Definition 3.
Let $S\!\subseteq\!\mathcal{D}$ be a sample. A neighbourhood $\mathcal{N}\!\subseteq\!S$ is called $S$-objective if $\mathbf{x}_{1},\mathbf{x}_{2}\!\in\!\mathcal{N}$ and there is an $\mathbf{x}_{3}\!\in S$ that is a convex combination of $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, then $\mathbf{x}_{3}\!\in\!\mathcal{N}$.
The next proposition explains that $S$-objectivity is always satisfied when we use an $\epsilon$-neighbourhood $\mathcal{N}_{\mathbf{x}}=\{\mathbf{x}^{\prime}\in S\ \mid\ \|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\epsilon\}$ of a point $\mathbf{x}$, that can be defined w.r.t. any common distance measure, including Euclidean, Manhattan and the Hamming distance and their weighted variants.
Proposition 1.
If $\mathcal{N}_{\mathbf{x}}$ is the $\epsilon$-neighbourhood of a point $\mathbf{x}\in S$ with respect to the distance induced by a seminorm [^2] A seminorm is a function $\|.\|:\mathcal{D}\rightarrow\mathcal{D}$ that satisfies sub-additivity (triangle inequality) and absolute homogeneity. $\|.\|$, then $\mathcal{N}$ is $S$-objective.
For the remaining properties, we introduce some notation.
Definition 4.
Given a finite neighbourhood $\mathcal{N}$ and a partial feature assignment $\mathbf{v}$ to a sequence of variables $\mathbf{V}$, the local probability of $\mathbf{v}$ in $\mathcal{N}$ is $P_{\mathcal{N}}(\mathbf{v})\!=\!\frac{|\{\mathbf{x}\in\mathcal{N}\ \mid\ (\mathbf{x}|_{\mathbf{V}})=\mathbf{v}\}|}{|\mathcal{N}|}$. The local success probability for $\mathbf{v}$ is defined as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid\mathbf{v})=\frac{|\{\mathbf{x}\in\mathcal{N}\ \mid\ \mathbf{x}|_{\mathbf{V}}=\mathbf{v},c(\mathbf{x})=1\}|}{|\{\mathbf{x}\in\mathcal{N}\ \mid\ \mathbf{x}|_{\mathbf{V}}=\mathbf{v}\}|}$.
Ideally, our neighbourhoods contain good representations of both the protected and non-protected groups. If it contained only one protected individual, one may wonder about other protected individuals. For example, even for a strongly biased model, we may be able to find one protected individual that is treated similar to non-protected ones even though the model strongly discriminates against the non-protected individuals in a majority of cases. Similarly, if it contained only one non-protected individual, this may simply be an outlier. Since protected groups are often underrepresented, we should not expect a balanced neighbourhood. However, if almost all elements in a neighbourhood belong to one group, the neighbourhood may not provide very strong evidence. To quantify neighbourhood diversity, we can use entropy:
Definition 5.
The entropy $H_{\mathcal{N}}(X_{p})$ of the marginal distribution of the protected feature is defined as
$H_{\mathcal{N}}(X_{p})=-P_{\mathcal{N}}(X_{p}=g)\cdot\log P_{\mathcal{N}}(X_{p}=g)$
$-P_{\mathcal{N}}(X_{p}\neq g)\cdot\log P_{\mathcal{N}}(X_{p}\neq g).$
The entropy takes its maximum $1$ when $\mathcal{N}$ is maximally diverse ($P_{\mathcal{N}}(X_{p}\!=\!g)\!\!=\!\!P_{\mathcal{N}}(X_{p}\!\neq\!g)\!=\!0.5)$, its minimum 0 when it is minimally so ($P_{\mathcal{N}}(X_{p}\!=\!g)\!\!=\!\!1$ or $P_{\mathcal{N}}(X_{p}\!\neq\!g)\!\!=\!\!1$), and is monotonically decreasing in between.
Definition 6.
Let $\epsilon\in[0,1]$. A neighbourhood $\mathcal{N}\subseteq\mathcal{D}$ is called $\epsilon$-diverse if $H_{\mathcal{N}}(g)=\epsilon$.
Finally, we define that a neighbourhood is biased if the local conditional probability of the positive outcome is significantly larger for the non-protected than the protected group.
Definition 7.
A neighbourhood $\mathcal{N}\subseteq\mathcal{D}$ is called $\epsilon$-biased against $X_{p}=g$ if $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)-P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)=\epsilon$ and $\epsilon>0$.
5 Argument scheme for detecting local bias
Our method makes use of properties of neighbourhoods to drive argumentative debates about the presence of bias against $X_{p}=g$. To shape these debates, we define novel argument schemes with accompanying critical questions. In the next section, we will turn both into QBAFs.
Figure 2 gives our argument scheme for detecting local bias against the protected feature taking the value of interest ($X_{p}=g$). The critical questions address the overall quality (properties) of the neighbourhood: CQ1. Is $\mathcal{N}$ NN-significant? CQ2. Is $\mathcal{N}$ $S$-objective? CQ3. Is $\mathcal{N}$ $\epsilon$-diverse?
When the critical questions are instantiated they give attacks against the conclusion of the (instantiated) argument scheme.
Major premise Generally, if, in $\mathcal{N}$,
$X_{p}\neq g$ leads to a positive decision
and $X_{p}=g$ leads to a negative decision,
then $X_{p}=g$ is disadvantaged in $\mathcal{N}$
and $X_{p}\neq g$ is advantaged in $\mathcal{N}$
Minor premise In the case of the chosen $\mathcal{N}$,
$X_{p}=g$ leads to a negative decision.
Minor premise In the case of a chosen $\mathcal{N}$,
$X_{p}\neq g$ leads to a positive decision.
Conclusion Thus, $X_{p}=g$ is disadvantaged in $\mathcal{N}$
and $X_{p}\neq g$ is advantaged in $\mathcal{N}$.
Figure 2: Argument scheme for neighbourhood $\mathcal{N}$. $X_{p}=g$ identifies the protected, potentially disadvantaged group.
We can also combine arguments drawn from different neighbourhoods, via the argument scheme in Figure 3. For simplicity, we will omit to consider any critical questions for this argument scheme, leaving this to future work.
Major premise If $X_{p}\!\!=\!\!g$ is disadvantaged in $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$,
then there is a bias against $X_{p}=g$.
Minor premise In the case of chosen $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$,
$X_{p}\!=\!g$ is disadvantaged in $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$.
Conclusion Thus, there is a bias against $X_{p}=g$.
Figure 3: Argument scheme for combining multiple neighbourhoods $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$.
6 Arguing about Bias
Our goal is to empower debates drawn from evidence for bias. To do so, first, to identify bias in a given neighbourhood $\mathcal{N}$, we introduce our core QBAF $Q_{\mathcal{N}}$ based on the argument scheme in Figure 2. Then, we extend $Q_{\mathcal{N}}$ with critical questions. Finally, we build our QBAF $\mathcal{Q}$ that allows to argue across neighbourhoods, based on the argument scheme from Figure 3. In this section, we assume a modular gradual semantics $\sigma$ with aggregation function $\operatorname{agg}$ and influence function $\operatorname{infl}$. We also omit the QBAF subscript when applying $\sigma$.
Arguing about Bias in a single Neighbourhood.
The conclusion of the argument scheme in Figure 2 amounts to two arguments, $\mathrm{Disadv}_{g}$ and $\mathrm{Adv}_{g}$, that express that individuals with $X_{p}=g$ are, resp., disadvantaged and advantaged. Their base score is set to 0. The minor premises give rise to two further arguments $\mathrm{Pos}_{g}$ and $\mathrm{Pos}_{\neq g}$. Their base score corresponds to the group’s local success probability in the neighbourhood. The major premise informs about the connection between these arguments: $\mathrm{Disadv}_{g}$ is supported by $\mathrm{Pos}_{g}$ and attacked by $\mathrm{Pos}_{\neq g}$; $\mathrm{Adv}_{g}$ is attacked by $\mathrm{Pos}_{g}$ and supported by $\mathrm{Pos}_{\neq g}$. Formally:
Definition 8.
The local bias-QBAF for $g$ w.r.t. neighbourhood $\mathcal{N}$ is the QBAF $Q_{\mathcal{N}}=(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ with
$\mathcal{A}=\{\mathrm{Disadv}_{g},\mathrm{Adv}_{g},\mathrm{Pos}_{g},\mathrm{Pos}_{\neq g}\}$,
$\mathcal{R}^{-}=\{(\mathrm{Pos}_{g},\mathrm{Disadv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Adv}_{g})\}$,
$\mathcal{R}^{+}=\{(\mathrm{Pos}_{g},\mathrm{Adv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Disadv}_{g})\}$, and
$\tau(a)\!=\!\begin{cases}P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g),&a=\mathrm{Pos}_{\neq g}\\ P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g),&a=\mathrm{Pos}_{g}.\\ 0,&else.\end{cases}$
Figure 1 shows this QBAF (yellow, right box). Intuitively, the base score of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ is 0 as our default assumption is that individuals with $X_{p}\!=\!g$ are not disadvantaged/advantaged. The base scores of $\mathrm{Pos}_{g}$$\mathrm{Pos}_{\neq g}$ are the local success probabilities of the groups $X_{p}\!=\!g$/ $X_{p}\!\neq\!g$, and 0 if it is never the case, $1$ if it is always the case, and otherwise a value between, corresponding to the relative frequency in $\mathcal{N}$.
We show that our choice of QBAF is sensible when choosing the semantics appropriately. First, intuitively, if $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$ and $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ are equal, then there is neither reason to accept that individuals with $X_{p}=g$ are disadvantaged nor that they are advantaged. However, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes larger than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Disadv}_{g}$. Conversely, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes smaller than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Adv}_{g}$. We can guarantee this behaviour by choosing aggregation and influence functions with certain properties.
Proposition 2.
If $\operatorname{agg}$ and $\operatorname{infl}$ satisfy
  1. 1.
    monotonicity, then $\sigma(\mathrm{Adv}_{g})\!\!=\!\!0$, $\sigma(\mathrm{Disadv}_{g})\!\!=\!\!0$ or both,
  2. 2.
    balance, then if $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)=P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, then $\sigma(\mathrm{Adv}_{g})=\sigma(\mathrm{Disadv}_{g})=0$,
  3. 3.
    strict monotonicity, then if $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)>P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, then $\sigma(\mathrm{Disadv}_{g})>0$,
  4. 4.
    strict monotonicity, then if $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)<P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, then $\sigma(\mathrm{Adv}_{g})>0$.
Item 1 states that we always fully reject at least one of $\mathrm{Adv}_{g}$ and $\mathrm{Disadv}_{g}$. We fully reject both if both groups have the same local success probability (item 2). Items 3 and 4 explain that if the protected/non-protected group is at a disadvantage, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ will be above 0.
Second, we should expect that, as the inequality between the protected and non-protected groups increases, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ should change accordingly. To study this situation, we compare the strength values in two independent local bias-QBAFs, connecting again properties of aggregation and influence functions with the expected behaviour:
Proposition 3.
Let $\mathcal{Q}$ be a QBAF composed of two independent local bias-QBAFs $Q_{\mathcal{N}^{1}},Q_{\mathcal{N}^{2}}$If $\operatorname{agg}$ and $\operatorname{infl}$ satisfy
  1. 1.
    balance, then, if $P_{\mathcal{N}^{1}}(c(\mathbf{X})\!=\!1\mid X_{p}\!\neq\!g)=P_{\mathcal{N}^{2}}(c(\mathbf{X})\!=\!1\mid X_{p}\!\neq\!g)$ and $P_{\mathcal{N}^{1}}(c(\mathbf{X})\!=\!1\mid X_{p}\!=\!g)\!=\!P_{\mathcal{N}^{2}}(c(\mathbf{X})\!=\!1\mid X_{p}\!=\!g)$, we have $\sigma(\mathrm{Adv}_{g}^{1})=\sigma(\mathrm{Adv}_{g}^{2})$ and $\sigma(\mathrm{Disadv}_{g}^{1})=\sigma(\mathrm{Disadv}_{g}^{2})$;
  2. 2.
    monotonicity, then if $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}\neq g)\geq P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ and $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}=g)\leq P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}=g)$, we have $\sigma(\mathrm{Adv}_{g}^{1})\leq\sigma(\mathrm{Adv}_{g}^{2})$ and $\sigma(\mathrm{Disadv}_{g}^{1})\geq\sigma(\mathrm{Disadv}_{g}^{2})$;
  3. 3.
    strict monotonicity, then the guarantees of item 2 hold, and, if additionally $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}\neq g)>P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ or $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}=g)<P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}=g)$, then $\sigma(\mathrm{Adv}_{g}^{1})<\sigma(\mathrm{Adv}_{g}^{2})$ and $\sigma(\mathrm{Disadv}_{g}^{1})>\sigma(\mathrm{Disadv}_{g}^{2})$;
  4. 4.
    monotonicity, then if $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}\neq g)\leq P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ and $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}=g)\geq P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}=g)$, we have $\sigma(\mathrm{Adv}_{g}^{1})\geq\sigma(\mathrm{Adv}_{g}^{2})$ and $\sigma(\mathrm{Disadv}_{g}^{1})\leq\sigma(\mathrm{Disadv}_{g}^{2})$;
  5. 5.
    strict monotonicity, then the guarantees of item 4 hold, and, if additionally $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}\neq g)<P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ or $P_{\mathcal{N}^{1}}(c(\mathbf{X})=1\mid X_{p}=g)>P_{\mathcal{N}^{2}}(c(\mathbf{X})=1\mid X_{p}=g)$, then $\sigma(\mathrm{Adv}_{g}^{1})>\sigma(\mathrm{Adv}_{g}^{2})$ and $\sigma(\mathrm{Disadv}_{g}^{1})<\sigma(\mathrm{Disadv}_{g}^{2})$.
Item 1 states that if we find the same inequality in two neighbourhoods, we will also have the same strengths. Items 2, 3 deal with the case where the protected group is disadvantaged. Item 2 is a generalization of item 1 and guarantees that if the non-protected/protected group in $\mathcal{N}_{1}$ is at least/most as successful as the non-protected/protected group in $\mathcal{N}_{2}$, then the strength of $\mathrm{Adv}_{g}$/ $\mathrm{Disadv}_{g}$ in $\mathcal{N}_{1}$ will be at least/most as large as in $\mathcal{N}_{2}$. Item 3 guarantees that there will be a strict difference in the strength values if there is a strict difference between the probabilities. Items 4, 5 give symmetrical guarantees when the protected group is advantaged.
Third, by construction, we treat evidence for being advantaged and disadvantaged equally. That is, in two neighbourhoods with symmetric evidence, the strength of $\mathrm{Adv}_{g}$ in one will be the strength of $\mathrm{Disadv}_{g}$ in the other:
Proposition 4.
For $Q_{\mathcal{N}^{1}},Q_{\mathcal{N}^{2}}$ as in Prop. 3, if $P_{\mathcal{N}^{1}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)\!=\!P_{\mathcal{N}^{2}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ and $P_{\mathcal{N}^{1}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)\!=\!P_{\mathcal{N}^{2}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$, then $\sigma(\mathrm{Adv}_{g}^{1})\!=\!\sigma(\mathrm{Disadv}_{g}^{2})$ and $\sigma(\mathrm{Adv}_{g}^{1})\!=\!\sigma(\mathrm{Disadv}_{g}^{2})$.
Finally, under DF-QuAD, the strength of the $\mathrm{Adv}_{g}$ and $\mathrm{Disadv}_{g}$ arguments has a particularly natural meaning:
Proposition 5.
For $\sigma$ the DF-QuAD semantics, we have
  1. 1.
    $\sigma(\mathrm{Adv}_{g})\!=\!\max\{0,P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)\!-\!P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\mid X_{p}\!\neq\!g)\}$,
  2. 2.
    $\sigma(\mathrm{Disadv}_{g})=\max\{0,P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)-P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)\}$,
  3. 3.
    $\sigma(\mathrm{Disadv}_{g})=\epsilon$ iff $\mathcal{N}$ is $\epsilon$-biased against $X_{p}=g$.
Questioning Bias in a single neighbourhood.
What if we detected bias only because the neighbourhood was chosen poorly? Here, we address this issue with critical questions from Section 5. We instantiate these questions by means of arguments and attacks added to local bias-QBAFs. We introduce three novel arguments, one per question: ss(ignificance) for CQ1, oo(bjectivity) for CQ2 and dd(iversity) for CQ3. By $CQ=\{s,o,d\}$, we denote the set of our novel arguments.
Intuitively, if the answer to these questions is negative, i.e., there is legitimate doubt regarding the significance, diversity, or objectivity of $\mathcal{N}$, then our confidence in detecting bias within $\mathcal{N}$ should decrease. Instead, if the answer is positive, then we can trust the bias detection outcome in $\mathcal{N}$. We thus treat the new arguments as attackers of $\mathrm{Adv}_{g}$ and $\mathrm{Disadv}_{g}$.
The new arguments are unattacked and unsupported by other arguments, and thus their strength (for any $\sigma$) will be their base score. So, for $s$, the base score should be monotonically decreasing w.r.t. the size of $\mathcal{N}$ and eventually reach 0, so that small neighbourhoods provide strong attacks. For the $o$, since objectivity is a binary criterion, an indicator function is most natural. For $d$, the attack strength should be maximal when the population consists of only one group, and approach 0 as the distribution becomes uniform: this behaviour can be captured by entropy, as given in Section 4. However, since we cannot expect a uniform distribution for underrepresented groups, we should be able to rescale the entropy: our base score function for $d$ should thus be monotonically decreasing. Formally, critical questions lead to the QBAF below:
Definition 9.
Let $Q_{\mathcal{N}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ be a local bias-QBAF for $g$ w.r.t. $\mathcal{N}$. Then the local bias-QBAF with critical questions for $g$ w.r.t. $\mathcal{N}$ is $\mathcal{Q}^{c}_{\mathcal{N}}\!=\!(\mathcal{A}^{\prime},{\mathcal{R}^{-}}^{\prime},\mathcal{R}^{+},\tau^{\prime})$ with $\mathcal{A}^{\prime}\!=\!\mathcal{A}\!\cup\!CQ$, ${\mathcal{R}^{-}}^{\prime}\!=\!\mathcal{R}^{-}\!\cup\!\{(a,\mathrm{Adv}_{g}),(a,\mathrm{Disadv}_{g})\!\mid\!a\!\in\!CQ\}$ and $\tau^{\prime}$ is such that for each $a\not\in CQ$, $\tau^{\prime}(a)=\tau(a)$ and for each $a\in CQ$, $\tau^{\prime}=f^{a}$ for functions $f^{a}$ as follows:
  • $f^{s}\!:\!\mathbb{N}\!\rightarrow\![0,1]$ takes in input $|\mathcal{N}|$ and: 1. it is monotonically decreasing, 2. $f^{s}(1)\!=\!1$, and 3. $\lim_{x\!\rightarrow\!\alpha}f^{s}(x)\!=\!0$ for some significance threshold $\alpha\!\in\!\mathbb{N}\!\cup\!\{\infty\}$;
  • for $S\!\subseteq\!\mathcal{D}$, $f^{o}\!:\!2^{\mathcal{D}}\!\rightarrow\!\{0,1\}$ takes in input $\mathcal{N}$ and is such that $f^{o}(\mathcal{N})\!=\!1$ if $\mathcal{N}$ is $S$-objective, and $f^{o}(\mathcal{N})\!=\!0$ else;
  • $f^{d}\!:\![0,1]\!\rightarrow\![0,1]$ takes in input the entropy $H_{\mathcal{N}}(g)$ of $\mathcal{N}$ and: 1. it is monotonically decreasing, 2. $f^{d}(0)\!=\!1$ and 3. $f^{d}(\beta)\!=\!0$ for some entropy threshold $\beta\in(0,1]$.
We illustrate this QBAF in Figure 1 (in the blue, left box). In the experiments (Section 7), we use $f^{s}(x)=\frac{\max\{\alpha-x,0\}}{\alpha}$ ($\alpha\!<\!\infty$) and $f^{d}(x)\!=\!1\!-\!\max\{\frac{x}{\beta},1\}$.
Similarly to Section 6, properties of $\sigma$ guarantee that the new arguments behave as intended. We do not go into formal detail but observe that, if $\operatorname{agg}$ and $\operatorname{infl}$ satisfy 1. balance, then $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will remain unaffected by critical questions that do not apply (base score 0), and 2. monotonicity, then the strength of $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will decrease monotonically w.r.t. the strength (i.e. base score) of $s,o,d$.
Arguing about Bias across neighbourhoods.
We combine local bias-QBAFs with critical questions across neighbourhoods $\mathcal{N}_{1},\dots,\mathcal{N}_{m}$, using the argument scheme in Figure 3, as follows. The conclusion yields a single argument ($\mathrm{bias}_{g}$}). We make the following assumptions: 1. we have no prior information about biases of the classifier and 2. in the absence of information, we see the classifier as unbiased (in dubio per reo). We capture this by setting the base score of $\mathrm{bias}_{g}$} to 0, meaning that we have no reason to believe there is a bias against $X_{p}\!=\!g$ to start with. The minor premise formulates that $X_{p}\!=\!g$ is disadvantaged in the neighbourhoods. We instantiate the detected disadvantage with two arguments per neighbourhood ($\mathrm{Adv}_{g}^{i}$ and $\mathrm{Disadv}_{g}^{i}$ in $\mathcal{N}_{i}$), giving $2m$ arguments. We use the major premise to identify an attack/support from each $\mathrm{Adv}_{g}^{i}$$\mathrm{Disadv}_{g}^{i}$ against $\mathrm{bias}_{g}$}. Formally:
Definition 10.
Let $Q^{c}_{\mathcal{N}_{i}}\!=\!(\mathcal{A}_{i},{\mathcal{R}_{i}}^{-},{\mathcal{R}_{i}}^{+},\tau_{i})$ be the local bias-QBAF with critical questions for $g$ w.r.t. $\mathcal{N}_{i}\!\in\!\{\mathcal{N}_{1},\dots,\mathcal{N}_{m}\}$. Then, the global bias-QBAF for $g$ is the QBAF $\mathcal{Q}_{\mathbf{g}}=(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ with
$\mathcal{A}=\{{\mathrm{bias}_{g}}\}\cup\bigcup_{1\leq i\leq m}\mathcal{A}_{i}$;
$\mathcal{R}^{-}=\{(\mathrm{Adv}_{g}^{i},{\mathrm{bias}_{g}})\mid 1\leq i\leq m\}\cup\bigcup_{i=1}^{m}\mathcal{R}_{i}^{-}$;
$\mathcal{R}^{+}=\{(\mathrm{Disadv}_{g}^{i},{\mathrm{bias}_{g}})\mid 1\leq i\leq m\}\cup\bigcup_{i=1}^{m}\mathcal{R}_{i}^{+}$;
$\tau(a)=\tau_{i}(a)$ if $a\in\mathcal{A}_{i}$ and 0 if $a={\mathrm{bias}_{g}}$.
Figure 1 illustrates a global bias-QBAF, for $m=2$.
Again, properties of $\sigma$ can guarantee that local bias-QBAFs affect $\mathrm{bias}_{g}$ as expected. In particular, $\mathrm{bias}_{g}$ will remain unaffected by fully rejected local bias arguments (strength 0) and its strength will increase/decrease monotonically w.r.t. local $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ arguments. Formally:
Proposition 6.
For $\mathcal{Q}_{\mathbf{g}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ a global bias-QBAF, let $S_{d}\!\!=\!\!\{\mathrm{Disadv}_{g}^{i}\mid i\!\in\!\{1,\ldots,m\},\sigma(\mathrm{Disadv}_{g}^{i})\!>\!0\}$ and $S_{a}\!=\!\{\mathrm{Adv}_{g}^{i}\!\mid\!i\!\in\!\{1,\ldots,m\},\sigma(\mathrm{Adv}_{g}^{i})\!>\!0\}$. If $\operatorname{agg}$ and $\operatorname{infl}$ satisfy balance and monotonicity, then 1. $S_{d}\succeq S_{a}$ implies $\sigma({\mathrm{bias}_{g}})\geq 0$, and 2. $S_{a}\succeq S_{d}$ implies $\sigma({\mathrm{bias}_{g}})=0$.
Note that item 1 of Proposition 2 implies that $S_{d}\cap S_{a}=\emptyset$.
7 Experiments
We conduct experiments with: 1. synthetically biased classifiers (two hand-crafted and one resulting from training on the Adult Census Income (ACI) dataset (Becker and Kohavi 1996), also used for testing ); 2. trained classifiers (logistic regression, but our method applies to any classifier); 3. LLMs (ChatGPT-4o). For training (2) and for testing (2,3) we use the Bank Marketing (Moro, Rita, and Cortez 2014) and COMPAS (ProPublica 2016; Angwin et al. 2016a) datasets. The datasets we choose are commonly used in the fairness literature (Räz 2024; Fawkes et al. 2024; Ehyaei, Farnadi, and Samadi 2024). We consider also LLMs as they are prone to societal biases due to their exposure to human-generated data (Ma et al. 2025; Xiang 2024). In all experiments, we compare with the argumentative approach by (Waller, Rodrigues, and Cocarascu 2024) (IRB in short). We report results with quadratic energy (DF-QuAD gave similar results). We report results for 1000 randomly sampled individuals and their KNN-neighbourhoods (Cover and Hart 1967), drawn from the test sets of the chosen datasets.
Experiments with Synthetically Biased Models.
We developed three synthetic models to obtain an objective ground truth in the setting of the ACI dataset: two globally biased models (Global 1 and Global 2) and one locally biased model (Local 1). Global 1 and 2 are hand-crafted decision trees (see Figures 5, 6 in Appendix C): the former predicts a negative label for all female individuals and a positive label for all others; the latter predicts a negative label for all black female individuals. For Local 1, we trained a logistic regression model on the ACI training set. In contrast to the global models, Local 1 operates at a neighbourhood level, changing the prediction to the negative label for female individuals if their neighbourhood exhibits an $\epsilon$-bias against $gender=female$ in the originally trained model (see Figure 7 in Appendix C).
Bias in Single Neighbourhoods. Table 1 shows that our method consistently outperforms IRB with Global 2 and Local 1 (IRB yields all zero scores with Global 2, as it focuses on showing bias against individual features, rather than combinations thereof). For Global 1, the two approaches perform similarly. Overall, our method runs significantly faster due to a simpler QBAF with fewer nodes and relations.
Method Model K Accuracy Precision Recall F1-score Runtime (s)
Our Global 1 50 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.90 $\pm$ 0.02 0.95 $\pm$ 0.01 3.87 $\pm$ 0.03
IRB 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.90 $\pm$ 0.01 0.95 $\pm$ 0.02 28.22 $\pm$ 0.04
Our Global 1 100 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 3.88 $\pm$ 0.02
IRB 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 48.47 $\pm$ 0.05
Our Global 1 200 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 3.82 $\pm$ 0.02
IRB 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 97.81 $\pm$ 0.04
Our Global 2 50 0.96 $\pm$ 0.02 1.00 $\pm$ 0.01 0.96 $\pm$ 0.02 0.98 $\pm$ 0.01 0.34 $\pm$ 0.02
IRB 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.28 $\pm$ 0.03
Our Global 2 100 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.35 $\pm$ 0.01
IRB 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 8.99 $\pm$ 0.04
Our Global 2 200 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.38 $\pm$ 0.01
IRB 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 17.44 $\pm$ 0.05
Our Local 1 50 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 5.00 $\pm$ 0.05
IRB 0.81 $\pm$ 0.02 1.00 $\pm$ 0.00 0.44 $\pm$ 0.02 0.61 $\pm$ 0.02 30.66 $\pm$ 0.05
Our Local 1 100 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 4.96 $\pm$ 0.04
IRB 0.74 $\pm$ 0.02 1.00 $\pm$ 0.00 0.31 $\pm$ 0.02 0.48 $\pm$ 0.02 49.45 $\pm$ 0.05
Our Local 1 200 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 5.07 $\pm$ 0.03
IRB 0.70 $\pm$ 0.01 1.00 $\pm$ 0.00 0.22 $\pm$ 0.01 0.36 $\pm$ 0.01 97.16 $\pm$ 0.04
Table 1: The performance of our method (using $Q^{c}_{\mathcal{N}_{i}}$) versus IRB, with single neighbourhoods ($\mathcal{N}_{i}$) of different sizes (K) for synthetically biased classifiers. Best results in bold.
Method Model Accuracy Precision Recall F1-score
Our Global 1 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
IRB 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
Our Global 2 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
IRB 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Our Local 1 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
IRB 0.70 $\pm$ 0.02 1.00 $\pm$ 0.01 0.22 $\pm$ 0.02 0.36 $\pm$ 0.01
Table 2: The performance of our method (using $\mathcal{Q}_{\mathbf{g}}$ with $\mathcal{N}_{1},\mathcal{N}_{2},\mathcal{N}_{3}$ of sizes K=50,100,200, respectively) versus IRB (with the largest neighbourhood $\mathcal{N}_{3}$) for synthetically biased classifiers. Best results in bold.
Bias Across Neighbourhoods. We consider multipleneighbourhoods of varying sizes (50, 100, 200). [^3] We run experiments with other sizes, ranging from $K=10$ to $K=200$ in intervals of 10. See Appendix C.2. Table 2 shows that our method achieves perfect performance for all classifiers. IRB fails for Global 2 for the same reasons as before.
Experiments with Trained Models.
We trained logistic regression models on Bank marketing and COMPAS. Here, we no longer have ground truth labels for bias and thus just compare the number of identified biases.
Feature value Logistic Regression
IRB Our
race
African-American 2 77
Caucasian 1 1
Hispanic 26 13
Asian 3 0
Native American 3 0
Other 11 0
sex
Female 1 0
(a) COMPAS dataset
Feature value Logistic Regression
IRB Our
age
MidAge 0 0
YoungOrOld 20 78
marital
Married 0 4
Single 4 11
Divorced 19 10
(b) Bank Marketing dataset
Table 3: Count of feature values being biased against in queried individuals using $\mathcal{Q}_{\mathbf{g}}$ with $\mathcal{N}_{1},\mathcal{N}_{2},\mathcal{N}_{3}$ of sizes K=50,100,200, resp., and IRB (with $\mathcal{N}_{3}$).
Table 3(a) gives results drawn from COMPAS, showing that our method identifies substantially more biased cases against African-American individuals (77 vs. 2), in line with the literature (Angwin et al. 2016b; Khademi and Honavar 2020), while detecting fewer or no biased cases for the other groups compared to IRB.Table 3(b) gives results for individuals identified as biased against specific age and marital status groups in the Bank Marketing dataset. In particular, for “YoungOrOld” age (younger than 25 years of age or older than 60 years of age), “Married”, and “Single” marital features, our method detects more biased cases than IRB. We show examples in Appendix C.
Experiments with LLMs.
We examine the widely used ChatGPT-4o. We followed the same protocol as in (Liu et al. 2024), prompting the model to predict the class label based on the individuals’ feature values. Table 4 shows the results of our experiments. Since the model is not fine-tuned on COMPAS or Bank Marketing, the results reveal the presence of inherent biases within the model itself. In Table 4(a), our method identifies a significantly higher number of biased instances against African Americans (129 cases) compared to zero by IRB. Similarly, the model demonstrates a higher degree of bias against females (6 instances). In Table 4(b), our approach also uncovers bias which IRB fails to detect. These results suggest that our method may be more sensitive in uncovering latent bias patterns that IRB overlooks.
Feature value LLM (GPT4)
IRB Our
race
African-American 0 129
Caucasian 0 2
Hispanic 8 1
Asian 2 0
Native American 3 0
Other 2 1
sex
Female 0 6
(a) COMPAS dataset
Feature value LLM (GPT4)
IRB Our
age
MidAge 0 2
YoungOrOld 0 14
marital
Married 0 13
Single 0 18
Divorced 0 0
(b) Bank Marketing dataset
Table 4: Count of feature values being biased against in queried individuals using $\mathcal{Q}_{\mathbf{g}}$ with $\mathcal{N}_{1},\mathcal{N}_{2},\mathcal{N}_{3}$ of sizes K=50,100,200, resp., versus IRB (with $\mathcal{N}_{3}$).
8 Argumentative Debates
Our QBAFs can be used as debate-based explanations. The debates are naturally argumentative, by presenting evidence for and against bias, and for and against this evidence. These debates may take place in the “mind” of a single agent, taking two opposing role, of a proponent arguing that there is a bias and an opponent arguing that there is no such bias, in the spirit of dispute trees and dispute derivations (Thang, Dung, and Hung 2009; Cyras et al. 2017). As an example, for the QBAF in Figure 1, the proponent may state the top most argument in Figure 4 (left) and the opponent may then continue by stating: “But some individuals amongst the 100 are of African-American race and do get parole!” This behaviour can be obtained with conversational templates as in (Rago et al. 2025). Also, the QBAFs can be used as templates to drive debates amongst two or more agents. To do so, roles (for and against bias) or specific neighbourhoods could be associated with agents, with each using critical question arguments to criticise the neighbourhoods chosen by the others. Finally, the QBAFs can be used as the backbone for debates between agents and humans (as in Figure 4). When debates arise within or amongst agents, they could be generated and presented as explanations to users in one go, or incrementally, based on the user’s prompts (as in Figure 4). This requires the ability of agents to match the prompts with specific parts of the underpinning QBAF, and generate replies based on relevant parts thereof. We leave this to future work.
Figure 4: Example of an Argumentative Debate (the debate in the yellow/blue boxes stems from the ones in Figure 1).
9 Conclusion
We presented a novel model-agnostic approach for bias detection for analysing both local and global biases through argument schemes and quantitative bipolar argumentation. Our approach addresses a critical need for algorithmic fairness, by integrating transparency into bias detection, while also exhibiting desirable properties, linked to established properties for quantitative bipolar argumentation, and performance advantages over an argumentative baseline.
We discussed how our approach may lead to argumentative debates: future work will focus on the realisation thereof within and across agents, as well as between agents and humans, including with user studies.

### v2

@@
-Argumentative Debates for Transparent Bias Detection [Technical Report]
+Argumentative Debates for Transparent Bias Detection
 Abstract
-As the use of AI in society grows, addressing emerging biases is essential to prevent systematic disadvantages against specific groups.
-Several fairness notions/algorithmic methods have been proposed, but, with very few exceptions, these tend to ignore transparency.
+As the use of AI in society grows, addressing emerging biases is essential to prevent systematic discrimination.
+Several bias detection methods have been proposed, but, with few exceptions, these tend to ignore transparency.
 Instead, interpretability and explainability are core requirements for algorithmic fairness, even more so than for other algorithmic solutions, given the human-oriented nature of fairness.
-In this paper, we contribute a novel transparent method for bias detection to empower debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods.
-Our method builds upon techniques from (formal and computational) argumentation, whereby debates result from arguing about biases within and across neighbourhoods.
-We evaluate our approach experimentally and demonstrate its strengths in performance against an argumentative baseline.
+We present ABIDE (Argumentative BIas detection by DEbate), a novel framework that structures bias detection transparently as debate, guided by an underlying argument graph as understood in (formal and computational) argumentation.
+The arguments are about the success chances of groups in local neighbourhoods and the significance of these neighbourhoods.
+We evaluate ABIDE experimentally and demonstrate its strengths in performance against an argumentative baseline.
 1 Introduction
-Figure 1: A QBAF generated by our approach for the COMPAS dataset (with protected feature $X_{p}\!=\!race$ and protected group g=g\!= African-American) for two neighbourhoods of size $K=10,100$ (critical question arguments with zero strength omitted).
-Dashed/solid edges are supports/atttacks.
-Green/red edges are supports/attacks (from arguments) with nonzero strength, and black edges are (from arguments) with zero strength.
-Edge width reflects the strength of originating arguments.
-(Strengths below/above nodes.
-Details in Section 6.)
 As the use of AI in society grows, addressing its potential unfairness is becoming increasingly crucial.
-For example, an unfair AI model used to support decision-making in healthcare, discriminating against a certain sub-population, would be extremely harmful.
+For example, an unfair AI model that supports decision-making in healthcare, discriminating against a certain sub-population, would be extremely harmful.
 In particular, it is essential to address any potential biases against specific groups/individuals in data and AI models trained thereupon (Mehrabi et al.
@@
 2017) defines fairness as an equal probability of being classified with the desirable (positive) label in different groups.
-However, existing solutions to identify and rectify unfairness tend to focus on optimising fairness and ignore the need for transparency for the roots of (un)fairness (Caton and Haas 2024).
-Instead, transparency, afforded by interpretability and explainability, is crucial for the detection and mitigation of bias, as is the case for any algorithmic solution, but even more so when bias is concerned, given that fairness is inherently human-oriented (Waller et al.
-2024).
-While considerable efforts have been made towards explainability of machine learning models (Dejl et al.
-2025; Ayoobi et al.
-2019, 2021, 2022, 2023a, 2023b, 2025; Leofante et al.
-2024), algorithmic fairness has received limited attention (exceptions include (Grabowicz, Perello, and Mishra 2022; Waller, Rodrigues, and Cocarascu 2024)).
-We propose a novel transparent method to detect bias through argumentative debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods.
+Existing solutions to identify unfairness tend to focus on optimising fairness and ignore the need for transparency for the roots of (un)fairness (Caton and Haas 2024).
+However, transparency, afforded by interpretability and explainability, is crucial for the trustworthy detection and mitigation of bias.
+While considerable efforts have been made towards explainability of machine learning models, e.g.
+as in (Ayoobi et al.
+2023b; Dejl et al.
+2025b), algorithmic fairness has received limited attention (exceptions include (Grabowicz, Perello, and Mishra 2022; Waller, Rodrigues, and Cocarascu 2024)).
+We propose ABIDE (Argumentative Bias Detection), a novel transparent method to detect bias through argumentative debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods.
 Debates recently emerged as a powerful mechanism to address various challenges in contemporary AI, such as AI safety (Brown-Cohen, Irving, and Piliouras 2024), and to improve LLM performance (Khan et al.
@@
 Thus, while these debates provide a justification for the outcomes, these outcomes are not faithfully explained by the debates.
-Our novel method for bias detection is designed to guide debates about the presence of bias against individuals.
+Figure 1: A QBAF generated by ABIDE for the COMPAS dataset (with protected feature $X_{p}\!=\!race$ and protected group g=g\!= African-American) for two neighbourhoods of size $K=10,100$ (critical question arguments with zero strength omitted).
+Dashed/solid edges are supports/atttacks.
+Green/red edges are supports/attacks (from arguments) with nonzero strength, and black edges are (from arguments) with zero strength.
+Edge width reflects the strength of originating arguments.
+(Strengths below/above nodes.
+Details in Section 6.)
+ABIDE is designed to guide debates about the presence of bias.
 To structure debates, we rely upon argument schemes (with critical questions) (Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017), from formal argumentation.
@@
     We show how our approach can empower debate-based bias detection in human-agent and multi-agent scenarios.
-The proofs of all results in the paper can be found in Appendix B. We will release the code on acceptance.
+The proofs of all results and additional material can be found in the accompanying (Ayoobi et al. 2025a). The code is available at https://github.com/hamed-ayoobi/ABIDE.
 2 Related Work
 Fairness in Machine Learning.
-Various methods for identifying, measuring and mitigating against bias have been proposed (Mehrabi et al.
+Various methods for identifying, measuring and mitigating bias have been proposed (Mehrabi et al.
 2022; Caton and Haas 2024), but they often neglect transparency (Waller et al.
 2024).
-One exception is (Grabowicz, Perello, and Mishra 2022), which proposed explanations to unearth and rectify bias, but relies upon feature importance, rather than debates, for transparency.
+One exception is (Grabowicz, Perello, and Mishra 2022), which uses explanations to unearth and rectify bias, but relies on feature importance, rather than debates, for transparency.
 Bias can be defined in different ways, e.g.
@@
 2017).
-Our approach is based on a local notion of statistical parity.
+ABIDE is based on a local notion of statistical parity.
 That is, we identify neighbourhoods where groups are locally advantaged/disadvantaged with respect to success probability.
@@
 Argumentation and Bias.
-(Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017) propose an argument scheme to capture arguments from bias111 //www.rationaleonline.com/map/9gy9gd/argument-from-bias, allowing to draw the conclusion that an arguer is unlikely to have taken both sides of an issue into account if the arguer is biased, subject to addressing critical questions regarding (i) evidence for the arguer’s bias and (ii) the need to take multiple sides for that issue.
+(Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017) propose an argument scheme to capture arguments from bias111www.rationaleonline.com/map/9gy9gd/argument-from-bias, allowing to draw the conclusion that an arguer is unlikely to have taken both sides of an issue into account if the arguer is biased, subject to addressing critical questions regarding (i) evidence for the arguer’s bias and (ii) the need to take multiple sides for that issue.
 Our new argument scheme below is orthogonal to this and can be used in combination with it, by addressing the first critical question using properties of neighbourhoods to collect evidence for the bias of the arguer (a classifier).
@@
 Argumentation and Debate.
+Our approach follows the general idea of using argumentation towards explainability, e.g.
+as in (Ayoobi et al.
+2019, 2021, 2022, 2023a, 2025b; Dejl et al.
+2025a).
 Several have pointed to the connection between argumentation and debate, e.g.
@@
 2025) for product recommendation.
-While we can leverage on these works towards debate generation, our approach is orthogonal.
+While we can leverage on these works towards debate generation, our approach is orthogonal to them.
 3 Preliminaries
@@
 For $\mathcal{D}\!=\!\bigtimes_{i=1}^{k}D_{i}$, vectors $\mathbf{x}\in\mathcal{D}$ are the inputs of the classification problem.
-We consider (binary) classifiers $\operatorname{c}:\mathcal{D}\rightarrow\mathcal{C}\!=\!\{0,1\}$.
-We regard $1$ as a desirable outcome (e.g., acceptance of a loan application) and 0 as a negative outcome.
+We consider (binary) classifiers $\operatorname{c}:\mathcal{D}\rightarrow\mathcal{C}\!=\!\{0,1\}$, and regard $1$ as a desirable (e.g., acceptance of a loan application) and 0 as a negative outcome.
 Notation 1.
@@
 QBAF Properties.
-QBAFs are often compared based on properties (Leite and Martins 2011; Amgoud and Ben-Naim 2016, 2017; Baroni, Rago, and Toni 2018), many of which are tied to properties of aggregation and influence functions (Mossakowski and Neuhaus 2018).
-We will use properties of (strict) monotonicity and balance of these functions from (Potyka and Booth 2024b) and recap them in Appendix A.
+QBAFs are often compared on properties (Leite and Martins 2011; Amgoud and Ben-Naim 2016, 2017; Baroni, Rago, and Toni 2018), many of which are tied to properties of aggregation and influence functions (Mossakowski and Neuhaus 2018).
+We will use properties of (strict) monotonicity and balance of these functions from (Potyka and Booth 2024b) and recap them in (Ayoobi et al.
+2025a).
 As shown in (Potyka and Booth 2024a), product and sum aggregation satisfy balance and monotonicity, sum satisfies strict monotonicity; the influence functions of DF-QuAD and quadratic energy satisfy all properties.
 4 Neighbourhoods and their Properties
-The goal of this section is to make the idea of local bias precise.
+In this section, we will formalize the idea of local bias.
 Intuitively, a local bias against a reference individual exists if other similar individuals are advantaged by the classifier.
-Formally, these similar individuals can be given by a neighbourhood of individuals surrounding the reference one.
+Formally, these similar individuals can be represented by a neighbourhood surrounding the reference individual.
 Definition 1.
@@
 To prevent this, one reasonable assumption is that when we pick two individuals from a sample, we cannot leave out individuals between them.
-Mathematically, betweenness can be described by convexity:
+Mathematically, betweenness can be described by convexity.
 Definition 3.
@@
 A neighbourhood $\mathcal{N}\!\subseteq\!S$ is called $S$-objective if $\mathbf{x}_{1},\mathbf{x}_{2}\!\in\!\mathcal{N}$ and there is an $\mathbf{x}_{3}\!\in S$ that is a convex combination of $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, then $\mathbf{x}_{3}\!\in\!\mathcal{N}$.
-The next proposition explains that $S$-objectivity is always satisfied when we use an $\epsilon$-neighbourhood $\mathcal{N}_{\mathbf{x}}=\{\mathbf{x}^{\prime}\in S\ \mid\ \|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\epsilon\}$ of a point $\mathbf{x}$, that can be defined w.r.t.
-any common distance measure, including Euclidean, Manhattan and the Hamming distance and their weighted variants.
+Formally, an input $\mathbf{x}\in\mathbb{R}^{n}$ is a convex combination of $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{n}$ if $\mathbf{x}=\lambda\cdot\mathbf{x}_{1}+(1-\lambda)\cdot\mathbf{x}_{2}$ for some $\lambda\in[0,1]$.
+Discrete ordinal features such as Boolean ({0,1}) or qualitative descriptions like small (0), medium (1), large (2) can be mapped to integers to match the definition.
+For nominal features such as color or marital status, we demand that they are equal in $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ (and thus in $\mathbf{x}$).
+The next proposition explains that $S$-objectivity is always satisfied when we use an $\epsilon$-neighbourhood $\mathcal{N}_{\mathbf{x}}=\{\mathbf{x}^{\prime}\in S\ \mid\ \|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\epsilon\}$ of a point $\mathbf{x}$.
+It can be defined w.r.t.
+any common distance measuresuch as Euclidean, Manhattan and the Hamming distance and their weighted variants.
 Proposition 1.
@@
 Ideally, our neighbourhoods contain good representations of both the protected and non-protected groups.
-If it contained only one protected individual, one may wonder about other protected individuals.
-For example, even for a strongly biased model, we may be able to find one protected individual that is treated similar to non-protected ones even though the model strongly discriminates against the non-protected individuals in a majority of cases.
+For example, even for a strongly biased model, we may be able to find one protected individual that is treated similar to non-protected ones even though the model strongly discriminates against the non-protected individuals in most cases.
 Similarly, if it contained only one non-protected individual, this may simply be an outlier.
-Since protected groups are often underrepresented, we should not expect a balanced neighbourhood.
+Since protected groups are often underrepresented, we should not expect balanced neighbourhoods.
 However, if almost all elements in a neighbourhood belong to one group, the neighbourhood may not provide very strong evidence.
-To quantify neighbourhood diversity, we can use entropy:
+To quantify neighbourhood diversity, we use entropy.
 Definition 5.
-The entropy $H_{\mathcal{N}}(X_{p})$ of the marginal distribution of the protected feature is defined as
-$H_{\mathcal{N}}(X_{p})=-P_{\mathcal{N}}(X_{p}=g)\cdot\log P_{\mathcal{N}}(X_{p}=g)$
-$-P_{\mathcal{N}}(X_{p}\neq g)\cdot\log P_{\mathcal{N}}(X_{p}\neq g).$
+The entropy $H_{\mathcal{N}}(X_{p})$ of the protected feature $X_{p}$ in $\mathcal{N}$ is $H_{\mathcal{N}}(X_{p})\!=\!-P_{\mathcal{N}}(X_{p}\!=\!g)\cdot\log P_{\mathcal{N}}(X_{p}\!=\!g)-P_{\mathcal{N}}(X_{p}\neq g)\cdot\log P_{\mathcal{N}}(X_{p}\neq g).$
 The entropy takes its maximum $1$ when $\mathcal{N}$ is maximally diverse ($P_{\mathcal{N}}(X_{p}\!=\!g)\!\!=\!\!P_{\mathcal{N}}(X_{p}\!\neq\!g)\!=\!0.5)$, its minimum 0 when it is minimally so ($P_{\mathcal{N}}(X_{p}\!=\!g)\!\!=\!\!1$ or $P_{\mathcal{N}}(X_{p}\!\neq\!g)\!\!=\!\!1$), and is monotonically decreasing in between.
@@
 A neighbourhood $\mathcal{N}\subseteq\mathcal{D}$ is called $\epsilon$-biased against $X_{p}=g$ if $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)-P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)=\epsilon$ and $\epsilon>0$.
-5 Argument scheme for detecting local bias
-Our method makes use of properties of neighbourhoods to drive argumentative debates about the presence of bias against $X_{p}=g$.
+5 Argument Schemes for Detecting Local Bias
+ABIDE makes use of properties of neighbourhoods to drive argumentative debates about the presence of bias against individuals with $X_{p}=g$.
 To shape these debates, we define novel argument schemes with accompanying critical questions.
@@
 Figure 2: Argument scheme for neighbourhood $\mathcal{N}$. $X_{p}=g$ identifies the protected, potentially disadvantaged group.
-We can also combine arguments drawn from different neighbourhoods, via the argument scheme in Figure 3.
-For simplicity, we will omit to consider any critical questions for this argument scheme, leaving this to future work.
+We can also combine arguments drawn from different neighbourhoods, via the argument scheme in Figure 3. For brevity, we omit to consider critical questions for this scheme.
 Major premise If $X_{p}\!\!=\!\!g$ is disadvantaged in $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$,
@@
 Figure 3: Argument scheme for combining multiple neighbourhoods $\mathcal{N}_{1},\ldots,\mathcal{N}_{m}$.
-6 Arguing about Bias
-Our goal is to empower debates drawn from evidence for bias.
-To do so, first, to identify bias in a given neighbourhood $\mathcal{N}$, we introduce our core QBAF $Q_{\mathcal{N}}$ based on the argument scheme in Figure 2.
-Then, we extend $Q_{\mathcal{N}}$ with critical questions.
-Finally, we build our QBAF $\mathcal{Q}$ that allows to argue across neighbourhoods, based on the argument scheme from Figure 3.
-In this section, we assume a modular gradual semantics $\sigma$ with aggregation function $\operatorname{agg}$ and influence function $\operatorname{infl}$.
-We also omit the QBAF subscript when applying $\sigma$.
+6 Arguing about Bias with ABIDE
+ABIDE empowers debates on bias, in three steps.
+  1. 1.
+    Build a local bias-QBAF $Q_{\mathcal{N}}$ (based on Figure 2).
+  2. 2.
+    Add critical questions (CQ1-CQ3).
+  3. 3.
+    Build a global bias-QBAF $\mathcal{Q}_{\mathbf{g}}$ (based on Figure 3).
+We assume a fixed modular gradual semantics $\sigma$ with aggregation function $\operatorname{agg}$ and influence function $\operatorname{infl}$ and omit the QBAF subscript when applying $\sigma$.
 Arguing about Bias in a single Neighbourhood.
-The conclusion of the argument scheme in Figure 2 amounts to two arguments, $\mathrm{Disadv}_{g}$ and $\mathrm{Adv}_{g}$, that express that individuals with $X_{p}=g$ are, resp., disadvantaged and advantaged.
+The conclusion of the argument scheme in Figure 2 amounts to two arguments, $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$, that express that individuals with $X_{p}=g$ are, resp., disadvantaged/advantaged.
 Their base score is set to 0.
-The minor premises give rise to two further arguments $\mathrm{Pos}_{g}$ and $\mathrm{Pos}_{\neq g}$.
+The minor premises give rise to two further arguments $\mathrm{Pos}_{=g}$ and $\mathrm{Pos}_{\neq g}$.
 Their base score corresponds to the group’s local success probability in the neighbourhood.
-The major premise informs about the connection between these arguments: $\mathrm{Disadv}_{g}$ is supported by $\mathrm{Pos}_{g}$ and attacked by $\mathrm{Pos}_{\neq g}$; $\mathrm{Adv}_{g}$ is attacked by $\mathrm{Pos}_{g}$ and supported by $\mathrm{Pos}_{\neq g}$.
-Formally:
+The major premise informs about the connection between these arguments: $\mathrm{Pos}_{=g}$ attacks/supports $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$, while $\mathrm{Pos}_{\neq g}$ supports/attacks $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$.
 Definition 8.
 The local bias-QBAF for $g$ w.r.t. neighbourhood $\mathcal{N}$ is the QBAF $Q_{\mathcal{N}}=(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ with
-$\mathcal{A}=\{\mathrm{Disadv}_{g},\mathrm{Adv}_{g},\mathrm{Pos}_{g},\mathrm{Pos}_{\neq g}\}$,
-$\mathcal{R}^{-}=\{(\mathrm{Pos}_{g},\mathrm{Disadv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Adv}_{g})\}$,
-$\mathcal{R}^{+}=\{(\mathrm{Pos}_{g},\mathrm{Adv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Disadv}_{g})\}$, and
-$\tau(a)\!=\!\begin{cases}P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g),&a=\mathrm{Pos}_{\neq g}\\ P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g),&a=\mathrm{Pos}_{g}.\\ 0,&else.\end{cases}$
+$\mathcal{A}=\{\mathrm{Disadv}_{g},\mathrm{Adv}_{g},\mathrm{Pos}_{=g},\mathrm{Pos}_{\neq g}\}$,
+$\mathcal{R}^{-}=\{(\mathrm{Pos}_{=g},\mathrm{Disadv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Adv}_{g})\}$,
+$\mathcal{R}^{+}=\{(\mathrm{Pos}_{=g},\mathrm{Adv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Disadv}_{g})\}$, and
+$\tau(a)\!=\!\begin{cases}P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g),&a=\mathrm{Pos}_{\neq g}\\ P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g),&a=\mathrm{Pos}_{=g}.\\ 0,&else.\end{cases}$
 Figure 1 shows this QBAF (yellow, right box).
 Intuitively, the base score of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ is 0 as our default assumption is that individuals with $X_{p}\!=\!g$ are not disadvantaged/advantaged.
-The base scores of $\mathrm{Pos}_{g}$$\mathrm{Pos}_{\neq g}$ are the local success probabilities of the groups $X_{p}\!=\!g$/ $X_{p}\!\neq\!g$, and 0 if it is never the case, $1$ if it is always the case, and otherwise a value between, corresponding to the relative frequency in $\mathcal{N}$.
+The base scores of $\mathrm{Pos}_{=g}$$\mathrm{Pos}_{\neq g}$ are the local success probabilities of the groups $X_{p}\!=\!g$/ $X_{p}\!\neq\!g$.
 We show that our choice of QBAF is sensible when choosing the semantics appropriately.
-First, intuitively, if $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$ and $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ are equal, then there is neither reason to accept that individuals with $X_{p}=g$ are disadvantaged nor that they are advantaged.
+If $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$ and $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ are equal, then there is neither reason to accept that individuals with $X_{p}=g$ are disadvantaged nor that they are advantaged.
 However, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes larger than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Disadv}_{g}$.
@@
   1. 1.
-    monotonicity, then $\sigma(\mathrm{Adv}_{g})\!\!=\!\!0$, $\sigma(\mathrm{Disadv}_{g})\!\!=\!\!0$ or both,
+    monotonicity, then $\sigma(\mathrm{Adv}_{g})\!\!=\!\!0$ or $\sigma(\mathrm{Disadv}_{g})\!\!=\!\!0$,
   2. 2.
@@
 Items 3 and 4 explain that if the protected/non-protected group is at a disadvantage, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ will be above 0.
-Second, we should expect that, as the inequality between the protected and non-protected groups increases, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ should change accordingly.
+As the inequality between the protected and non-protected groups increases, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ should change accordingly.
 To study this situation, we compare the strength values in two independent local bias-QBAFs, connecting again properties of aggregation and influence functions with the expected behaviour:
@@
 Items 4, 5 give symmetrical guarantees when the protected group is advantaged.
-Third, by construction, we treat evidence for being advantaged and disadvantaged equally.
+By design, we treat evidence for being advantaged and disadvantaged equally.
 That is, in two neighbourhoods with symmetric evidence, the strength of $\mathrm{Adv}_{g}$ in one will be the strength of $\mathrm{Disadv}_{g}$ in the other:
@@
 Questioning Bias in a single neighbourhood.
-What if we detected bias only because the neighbourhood was chosen poorly?
-Here, we address this issue with critical questions from Section 5.
-We instantiate these questions by means of arguments and attacks added to local bias-QBAFs.
-We introduce three novel arguments, one per question: ss(ignificance) for CQ1, oo(bjectivity) for CQ2 and dd(iversity) for CQ3.
+The significance of the evidence given by a neighbourhood can be debatable.
+To take account of this, we introduce additional arguments corresponding to the critical questions: ss(ignificance) for CQ1, oo(bjectivity) for CQ2 and dd(iversity) for CQ3.
 By $CQ=\{s,o,d\}$, we denote the set of our novel arguments.
@@
 the size of $\mathcal{N}$ and eventually reach 0, so that small neighbourhoods provide strong attacks.
-For the $o$, since objectivity is a binary criterion, an indicator function is most natural.
-For $d$, the attack strength should be maximal when the population consists of only one group, and approach 0 as the distribution becomes uniform: this behaviour can be captured by entropy, as given in Section 4.
+For $o$, since objectivity is a binary criterion, an indicator function is most natural.
+For $d$, the attack strength should be maximal when the population consists of only one group, and approach 0 as the distribution becomes uniform: this behaviour can be captured by the entropy, as given in Section 4.
 However, since we cannot expect a uniform distribution for underrepresented groups, we should be able to rescale the entropy: our base score function for $d$ should thus be monotonically decreasing.
@@
 Definition 9.
-Let $Q_{\mathcal{N}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ be a local bias-QBAF for $g$ w.r.t.
-$\mathcal{N}$.
-Then the local bias-QBAF with critical questions for $g$ w.r.t.
-$\mathcal{N}$ is $\mathcal{Q}^{c}_{\mathcal{N}}\!=\!(\mathcal{A}^{\prime},{\mathcal{R}^{-}}^{\prime},\mathcal{R}^{+},\tau^{\prime})$ with $\mathcal{A}^{\prime}\!=\!\mathcal{A}\!\cup\!CQ$, ${\mathcal{R}^{-}}^{\prime}\!=\!\mathcal{R}^{-}\!\cup\!\{(a,\mathrm{Adv}_{g}),(a,\mathrm{Disadv}_{g})\!\mid\!a\!\in\!CQ\}$ and $\tau^{\prime}$ is such that for each $a\not\in CQ$, $\tau^{\prime}(a)=\tau(a)$ and for each $a\in CQ$, $\tau^{\prime}=f^{a}$ for functions $f^{a}$ as follows:
-  • $f^{s}\!:\!\mathbb{N}\!\rightarrow\![0,1]$ takes in input $|\mathcal{N}|$ and: 1.
-it is monotonically decreasing, 2.
-$f^{s}(1)\!=\!1$, and 3.
-$\lim_{x\!\rightarrow\!\alpha}f^{s}(x)\!=\!0$ for some significance threshold $\alpha\!\in\!\mathbb{N}\!\cup\!\{\infty\}$;
-  • for $S\!\subseteq\!\mathcal{D}$, $f^{o}\!:\!2^{\mathcal{D}}\!\rightarrow\!\{0,1\}$ takes in input $\mathcal{N}$ and is such that $f^{o}(\mathcal{N})\!=\!1$ if $\mathcal{N}$ is $S$-objective, and $f^{o}(\mathcal{N})\!=\!0$ else;
-  • $f^{d}\!:\![0,1]\!\rightarrow\![0,1]$ takes in input the entropy $H_{\mathcal{N}}(g)$ of $\mathcal{N}$ and: 1.
-it is monotonically decreasing, 2.
-$f^{d}(0)\!=\!1$ and 3.
-$f^{d}(\beta)\!=\!0$ for some entropy threshold $\beta\in(0,1]$.
+Let $Q_{\mathcal{N}}\!=\!(\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau)$ be a local bias-QBAF.
+The corresponding local bias-QBAF with critical questions is $\mathcal{Q}^{c}_{\mathcal{N}}\!=\!(\mathcal{A}^{\prime},{\mathcal{R}^{-}}^{\prime},\mathcal{R}^{+},\tau^{\prime})$ with $\mathcal{A}^{\prime}\!=\!\mathcal{A}\!\cup\!CQ$, ${\mathcal{R}^{-}}^{\prime}\!=\!\mathcal{R}^{-}\!\cup\!\{(a,\mathrm{Adv}_{g}),(a,\mathrm{Disadv}_{g})\!\mid\!a\!\in\!CQ\}$ and $\tau^{\prime}$ is such that $\tau^{\prime}(a)=\tau(a)$ for all $a\not\in CQ$, and
+  • $\tau^{\prime}(s)=f^{s}(|\mathcal{N}|)$ for a monotonically decreasing function $f^{s}\!:\!\mathbb{N}\!\rightarrow\![0,1]$ such that $f^{s}(1)\!=\!1$, and $\lim_{x\!\rightarrow\!\alpha}f^{s}(x)\!=\!0$ for some significance threshold $\alpha\!\in\!\mathbb{N}\!\cup\!\{\infty\}$;
+  • $\tau^{\prime}(o)=f^{o}(\mathcal{N})$ where $f^{o}(\mathcal{N})\!=\!0$ if $\mathcal{N}$ is $S$-objective, and $f^{o}(\mathcal{N})\!=\!\gamma$ else, where $\gamma\in(0,1]$ is an objectivity weight;
+  • $\tau^{\prime}(d)=f^{d}(H_{\mathcal{N}}(g))$ where for $f^{d}\!:\![0,1]\!\rightarrow\![0,1]$ is monotonically decreasing and satisfies $f^{d}(0)\!=\!1$ and $f^{d}(\beta)\!=\!0$ for some entropy threshold $\beta\in(0,1]$.
 We illustrate this QBAF in Figure 1 (in the blue, left box).
@@
 Similarly to Section 6, properties of $\sigma$ guarantee that the new arguments behave as intended.
-We do not go into formal detail but observe that, if $\operatorname{agg}$ and $\operatorname{infl}$ satisfy 1.
-balance, then $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will remain unaffected by critical questions that do not apply (base score 0), and 2.
-monotonicity, then the strength of $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will decrease monotonically w.r.t.
+We do not go into formal detail but observe that, if $\operatorname{agg}$ and $\operatorname{infl}$ satisfy balance, then $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will remain unaffected by critical questions that do not apply (base score 0), and, if they satisfy monotonicity, then the strength of $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ will decrease monotonically w.r.t.
 the strength (i.e.
@@
 Arguing about Bias across neighbourhoods.
-We combine local bias-QBAFs with critical questions across neighbourhoods $\mathcal{N}_{1},\dots,\mathcal{N}_{m}$, using the argument scheme in Figure 3, as follows.
-The conclusion yields a single argument ($\mathrm{bias}_{g}$}).
-We make the following assumptions: 1.
-we have no prior information about biases of the classifier and 2.
-in the absence of information, we see the classifier as unbiased (in dubio per reo).
-We capture this by setting the base score of $\mathrm{bias}_{g}$} to 0, meaning that we have no reason to believe there is a bias against $X_{p}\!=\!g$ to start with.
-The minor premise formulates that $X_{p}\!=\!g$ is disadvantaged in the neighbourhoods.
-We instantiate the detected disadvantage with two arguments per neighbourhood ($\mathrm{Adv}_{g}^{i}$ and $\mathrm{Disadv}_{g}^{i}$ in $\mathcal{N}_{i}$), giving $2m$ arguments.
-We use the major premise to identify an attack/support from each $\mathrm{Adv}_{g}^{i}$$\mathrm{Disadv}_{g}^{i}$ against $\mathrm{bias}_{g}$}.
-Formally:
+Finally, we combine multiple local bias-QBAFs for different neighbourhoods $\mathcal{N}_{1},\dots,\mathcal{N}_{m}$, as per Figure 3.
+Their individual $\mathrm{Adv}_{g}/\mathrm{Disadv}_{g}$ arguments attack/support a global bias argument $\mathrm{bias}_{g}$}.
+We make the following assumptions: we have no prior information about biases of the classifier and, in the absence of information, we take that the classifier is unbiased.
+We capture this by setting the base score of $\mathrm{bias}_{g}$} to 0.
 Definition 10.
@@
 local $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ arguments.
-Formally:
 Proposition 6.
@@
 For training (2) and for testing (2,3) we use the Bank Marketing (Moro, Rita, and Cortez 2014) and COMPAS (ProPublica 2016; Angwin et al.
-2016a) datasets.
+2016) datasets.
 The datasets we choose are commonly used in the fairness literature (Räz 2024; Fawkes et al.
@@
 2025; Xiang 2024).
-In all experiments, we compare with the argumentative approach by (Waller, Rodrigues, and Cocarascu 2024) (IRB in short).
+In all experiments, we compare ABIDE with the argumentative approach by (Waller, Rodrigues, and Cocarascu 2024) (IRB in short).
 We report results with quadratic energy (DF-QuAD gave similar results).
 We report results for 1000 randomly sampled individuals and their KNN-neighbourhoods (Cover and Hart 1967), drawn from the test sets of the chosen datasets.
+A sensitivity analysis evaluating the model’s performance across varying parameters is provided in (Ayoobi et al.
+2025a).
 Experiments with Synthetically Biased Models.
 We developed three synthetic models to obtain an objective ground truth in the setting of the ACI dataset: two globally biased models (Global 1 and Global 2) and one locally biased model (Local 1).
-Global 1 and 2 are hand-crafted decision trees (see Figures 5, 6 in Appendix C): the former predicts a negative label for all female individuals and a positive label for all others; the latter predicts a negative label for all black female individuals.
+Global 1 and 2 are hand-crafted decision trees (see Figures 5, 6 in Appendix 3): the former predicts a negative label for all female individuals and a positive label for all others; the latter predicts a negative label for all black female individuals.
 For Local 1, we trained a logistic regression model on the ACI training set.
-In contrast to the global models, Local 1 operates at a neighbourhood level, changing the prediction to the negative label for female individuals if their neighbourhood exhibits an $\epsilon$-bias against $gender=female$ in the originally trained model (see Figure 7 in Appendix C).
+In contrast to the global models, Local 1 operates at a neighbourhood level, changing the prediction to the negative label for female individuals if their neighbourhood exhibits an $\epsilon$-bias against $gender=female$ in the originally trained model (see Figure 7 in (Ayoobi et al.
+2025a)).
 Bias in Single Neighbourhoods.
@@
 [^3] We run experiments with other sizes, ranging from $K=10$ to $K=200$ in intervals of 10.
-See Appendix C.2.
+See (Ayoobi et al.
+2025a) for details.
 Table 2 shows that our method achieves perfect performance for all classifiers.
@@
 2), in line with the literature (Angwin et al.
-2016b; Khademi and Honavar 2020), while detecting fewer or no biased cases for the other groups compared to IRB.Table 3(b) gives results for individuals identified as biased against specific age and marital status groups in the Bank Marketing dataset.
+2016; Khademi and Honavar 2020), while detecting fewer or no biased cases for the other groups compared to IRB.
+Table 3(b) gives results for individuals identified as biased against specific age and marital status groups in the Bank Marketing dataset.
 In particular, for “YoungOrOld” age (younger than 25 years of age or older than 60 years of age), “Married”, and “Single” marital features, our method detects more biased cases than IRB.
-We show examples in Appendix C.
+We show examples in (Ayoobi et al.
+2025a).
 Experiments with LLMs.
@@
 We leave this to future work.
-Figure 4: Example of an Argumentative Debate (the debate in the yellow/blue boxes stems from the ones in Figure 1).
+Figure 4: Example of an Argumentative Debate (the debates in the yellow/blue boxes stem from the corresponding boxes in Figure 1).
 9 Conclusion
-We presented a novel model-agnostic approach for bias detection for analysing both local and global biases through argument schemes and quantitative bipolar argumentation.
+We presented ABIDE, a novel model-agnostic approach for bias detection for analysing both local and global biases through argument schemes and quantitative bipolar argumentation.
 Our approach addresses a critical need for algorithmic fairness, by integrating transparency into bias detection, while also exhibiting desirable properties, linked to established properties for quantitative bipolar argumentation, and performance advantages over an argumentative baseline.
-We discussed how our approach may lead to argumentative debates: future work will focus on the realisation thereof within and across agents, as well as between agents and humans, including with user studies.
+We discussed how ABIDE can support debates about bias.
+Future work will focus on the realisation thereof within and across agents, as well as between agents and humans, including with user studies.
+It will be especially interesting to explore how humans can contribute to the debates, specifically by contesting existing arguments and adding new arguments to the debate, in the spirit of (Leofante et al.
+2024).
+Acknowledgements
+Ayoobi, Rapberger, and Toni were funded by the ERC under the EU’s Horizon 2020 research and innovation programme (ADIX, grant number 101020934).
+Toni was also funded by J.P.
+Morgan and by the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme.